# FreshBrew: A Benchmark for Evaluating AI Agents on Java Code Migration

### v1

@@ 1 Introduction
Moving forward is, however, painful. Oracle’s own migration manual cautions that: "every new Java SE release introduces some binary, source, and behavioural incompatibilities" [^2] https://docs.oracle.com/en/java/javase/11/migrate/index.html. Java libraries also evolve in breaking ways: Raemaekers et al. (2017) examined >22,000 Maven artifacts and observed that $\approx\frac{1}{3}$ of all releases introduce at least one breaking change, regardless of whether the version bump is major or minor.
The tasks of upgrading Java version and upgrading dependency versions are intertwined. Some newer dependency may no longer support older Java versions. For example, Spring Boot 3.0 is a cornerstone for thousands of enterprises and ships on a Java 17 baseline; upgrading the framework (e.g., for security fixes or Jakarta EE 10 support) is therefore impossible without first moving the application to JDK 17  [^3] https://spring.io/blog/2022/11/24/spring-boot-3-0-goes-ga. Conversely, some older dependencies do not support newer Java versions. For example, libraries like Netty, Mockito, and Hazelcast relied on sun.misc.Unsafe for performance-critical operations. The encapsulation of this internal class in Java 9 forced these libraries to be completely re-engineered  [^4] https://www.infoq.com/articles/java-9-sun-misc-unsafe/.
@@ 2.2 Benchmark Datasets for Repository-Level Code Migration
One notable exception is the concurrent work of MigrationBench Liu et al. (2025), which targets Java 8 to Java 17+ migrations at the repository level. While it represents a significant step toward realistic evaluation, the current benchmark design exhibits several critical limitation, particularly the absence of safeguards against reward hacking. Specifically, MigrationBench neither selects high-coverage repositories nor enforces the preservation of test coverage. As a result, agents may superficially “succeed” by deleting or commenting out failing tests, removing problematic modules, or modifying build configurations to suppress errors. These loopholes allow systems to artificially inflate test pass rates without preserving the original program semantics, thereby undermining the benchmark’s ability to assess correctness and maintainability accurately. Moreover, the work does not include experiments involving AI agents, nor does it explicitly address the issue of agent reward hacking.
In contrast, our benchmark selectively curates a dataset consisting of repositories with high test coverage and implements an evaluation protocol that explicitly enforces coverage preservation. This design mitigates opportunities for reward hacking and ensures that successful migrations reflect genuine semantic correctness. Additionally, we provide a comprehensive evaluation of state-of-the-art LLM-based agents under these stricter, more realistic conditions.
Furthermore, our evaluation protocol differs significantly. FreshBrew is designed to evaluate multi-tool AI agents that can interact with a file system, run build commands, and use web search to solve problems. In contrast, MigrationBench evaluates the direct outputs of LLMs without this agentic tool-use layer, which we argue is critical for assessing the practical utility of modern AI coding assistants.
@@ 3.2 Evaluation Protocol
  • Cost. We measure the total cost of using the LLM during a migration run. This metric directly correlates with overall latency. We measure the cost of utilizing each agent by calculating the expense based on the per-token input and output pricing for each LLM, as provided by the together.ai [^8] http://together.ai/ API.
@@ 4.1.1 Tool-Augmented agent
We conducted experiments with a CodeAct Wang et al. (2024) agent, as implemented by the smolagents Roucher et al. (2025) framework 999huggingface.co/docs/smolagents/en/reference/agents#smolagents.CodeAgent. To ensure comprehensive coverage, we evaluated a diverse subset of models, including open-weight models, enterprise-grade models, and specialized coding models.
@@ 4.1.2 Deterministic Baseline with OpenRewrite
To contextualize the performance of the AI agents, we established a baseline using OpenRewrite, a state-of-the-art deterministic refactoring tool. We evaluated its ability to perform the migration using the composite recipe java.migrate.UpgradeToJava21  [^10] https://docs.openrewrite.org/recipes/java/migrate/upgradetojava21. This recipe programmatically applies a series of fine-grained transformations, such as updating Maven compiler settings and replacing deprecated APIs, by operating on a Lossless Semantic Tree (LST) representation of the source code.
@@ 4.2 Experimental Results
As expected, migrating to JDK 21 proved to be a more challenging task, with all models exhibiting a drop in performance compared to the JDK 17 task. For instance, the top-performing model, Gemini 2.5 Flash, saw its success rate decrease from 52.3% on the JDK 17 task to 52.4% on the JDK 21 task. This trend highlights the increasing complexity and difficulty of migrating to newer Java versions. This observation is discussed in more detail in Section 4.3.2.
@@ Appendix B Prompts
{promptcode}
Upgrade the Maven project rooted at the current working dir to use JDK version. Ensure that the build and the tests pass. You can make source code and dependency changes. You cannot delete or comment-out tests. To do a final verification of your work before calling the ‘final_answer‘ tool, verify your work by making sure that ‘maven_verify‘ tool succeeds with the parameter ‘clean‘ set to ‘True‘, because this is how we are going to evaluate your success. DO NOT call maven_verify with ‘clean=True‘ before ensuring it passes without it, because its slow and outputs a lot of tokens.
@@
{promptcode}
You are an expert Java software engineer and researcher specializing in code migration and developer tool evaluation. Your task is to analyze the final n_steps ’thought-action’ steps from a failed attempt by an AI agent to migrate a Java 8 project to Java 17. Based on the provided trace, identify the primary technical reason for the failure. Do not simply state that the agent failed or ran out of steps. Pinpoint the specific build, dependency, or code-level issue that the agent was unable to resolve. Choose ONLY ONE of the following categories that best describes the failure: * Dependency Management Failure * Build Configuration Error * Java API Incompatibility * Agent Behavioral Failure * Root Cause Not in Final Steps * Unknown The agent’s final n_steps steps are as follows: — BEGIN TRACE — final_steps_trace — END TRACE — Provide your output in JSON format with two keys: "failure_category" and "reasoning". The reasoning should be a brief, one-sentence explanation supporting your choice.

### v2

@@ 1 Introduction
Moving forward is, however, painful. Oracle’s own migration manual cautions that: "every new Java SE release introduces some binary, source, and behavioural incompatibilities"222docs.oracle.com/en/java/javase/11/migrate/index.html. Java libraries also evolve in breaking ways: Raemaekers et al. (2017) examined >22,000 Maven artifacts and observed that $\approx\frac{1}{3}$ of all releases introduce at least one breaking change, regardless of whether the version bump is major or minor.
The tasks of upgrading Java version and upgrading dependency versions are intertwined. Some newer dependency may no longer support older Java versions. For example, Spring Boot 3.0 is a cornerstone for thousands of enterprises and ships on a Java 17 baseline; upgrading the framework (e.g., for security fixes or Jakarta EE 10 support) is therefore impossible without first moving the application to JDK 17 333spring.io/blog/2022/11/24/spring-boot-3-0-goes-ga. Conversely, some older dependencies do not support newer Java versions. For example, libraries like Netty, Mockito, and Hazelcast relied on sun.misc.Unsafe for performance-critical operations. The encapsulation of this internal class in Java 9 forced these libraries to be completely re-engineered 444www.infoq.com/articles/java-9-sun-misc-unsafe/.
@@ 2.2 Benchmark Datasets for Repository-Level Code Migration
One notable exception is the concurrent [^8] MigrationBench was first released publicly on arXiv in May 2025, while FreshBrew was developed independently and submitted in July 2025, with its arXiv preprint first posted in October 2025. work of MigrationBench (Liu et al., 2025), which introduces a repository-level benchmark for migrating Java 8 projects to JDK 17+. It represents a major step toward realistic large-scale evaluation, with a detailed protocol that checks build success, verifies test integrity, and distinguishes minimal vs. maximal migrations. MigrationBench also locates the last buildable Java 8 revision in each repository’s history, ensuring valid starting points for migration.
In contrast, FreshBrew targets evaluation challenges specific to agentic systems, where models actively manipulate files, execute builds, and use tools to perform migration tasks. Such agents are prone to reward hacking—appearing successful by deleting failing tests, removing problematic modules, or altering build settings to suppress errors. To detect and prevent these behaviors, FreshBrew (1) curates high-coverage repositories, (2) enforces test-coverage preservation, and (3) conducts experiments centered on multi-tool AI agents under this protocol. This experimental focus highlights the distinctive failure modes and reward-hacking patterns that arise in autonomous coding agents, complementing benchmarks like MigrationBench that evaluate non-agentic settings.
@@ 3.2 Evaluation Protocol
  • Cost. We measure the total cost of using the LLM during a migration run. This metric directly correlates with overall latency. We measure the cost of utilizing each agent by calculating the expense based on the per-token input and output pricing for each LLM, as provided by the together.ai999together.ai/ API.
@@ 4.1.1 Tool-Augmented agent
We conducted experiments with a CodeAct Wang et al. (2024) agent, as implemented by the smolagents Roucher et al. (2025) framework 101010huggingface.co/docs/smolagents/en/reference/agents#smolagents.CodeAgent. To ensure comprehensive coverage, we evaluated a diverse subset of models, including open-weight models, enterprise-grade models, and specialized coding models.
@@ 4.1.2 Deterministic Baseline with OpenRewrite
To contextualize the performance of the AI agents, we established a baseline using OpenRewrite, a state-of-the-art deterministic refactoring tool. We evaluated its ability to perform the migration using the composite recipe java.migrate.UpgradeToJava21 111111docs.openrewrite.org/recipes/java/migrate/upgradetojava21. This recipe programmatically applies a series of fine-grained transformations, such as updating Maven compiler settings and replacing deprecated APIs, by operating on a Lossless Semantic Tree (LST) representation of the source code.
@@ 4.2 Experimental Results
As expected, migrating to JDK 21 proved to be a more challenging task, with all models exhibiting a drop in performance compared to the JDK 17 task. For instance, the top-performing model, Gemini 2.5 Flash, saw its success rate decrease from 52.3% on the JDK 17 task to 49.8% on the JDK 21 task. This trend highlights the increasing complexity and difficulty of migrating to newer Java versions. This observation is discussed in more detail in Section 4.3.2.
@@ Appendix B Prompts
Upgrade the Maven project rooted at the current working dir to use JDK {version}. Ensure that the build and the tests pass. You can make source code and dependency changes. You cannot delete or comment-out tests. To do a final verification of your work before calling the final_answer tool, verify your work by making sure that maven_verify tool succeeds with the parameter clean set to True‘, because this is how we are going to evaluate your success. DO NOT call maven_verify with clean=True before ensuring it passes without it, because its slow and outputs a lot of tokens.
@@
You are an expert Java software engineer and researcher specializing in code migration and developer tool evaluation. Your task is to analyze the final {{n_steps}} thought-action steps from a failed attempt by an AI agent to migrate a Java 8 project to Java 17.
Based on the provided trace, identify the primary technical reason for the failure. Do not simply state that the agent failed or ran out of steps. Pinpoint the specific build, dependency, or code-level issue that the agent was unable to resolve.
Choose ONLY ONE of the following categories that best describes the failure:
* Dependency Management Failure
* Build Configuration Error
* Java API Incompatibility
* Agent Behavioral Failure
* Root Cause Not in Final Steps
* Unknown
The agents final {{n_steps}} steps are as follows:
--- BEGIN TRACE ---
{{final_steps_trace}}
--- END TRACE ---
Provide your output in JSON format with two keys: "failure_category" and "reasoning". The reasoning should be a brief, one-sentence explanation supporting your choice.